### v1

@@ Abstract
Considering the close connection between action recognition and human pose estimation, we design a Collaboratively Self-supervised Video Representation (CSVR) learning framework specific to action recognition by jointly considering generative pose prediction and discriminative context matching as pretext tasks. Specifically, our CSVR consists of three branches: a generative pose prediction branch, a discriminative context matching branch, and a video generating branch. Among them, the first one encodes dynamic motion feature by utilizing Conditional-GAN to predict the human poses of future frames, and the second branch extracts static context features by pulling the representations of clips and compressed key frames from the same video together while pushing apart the pairs from different videos. The third branch is designed to recover the current video frames and predict the future ones, for the purpose of collaboratively improving dynamic motion features and static context features. Extensive experiments demonstrate that our method achieves state-of-the-art performance on the UCF101 and HMDB51 datasets.
keywords:
@@
Video representation learning, Self-supervised learning, Action recognition, Human pose prediction
1 Introduction
@@
Human action recognition is a fundamental task in computer vision and has gained increasing attention in recent years. With a large number of labeled videos, the convolutional neural network (CNN) has achieved great success in human action recognition by learning powerful spatiotemporal representation features [32, 33, 81, 80]. However, the cost of collecting and tagging large amounts of videos is enormous. Fortunately, there is a huge amount of unlabeled videos on the Internet. A common practice is to use a large amount of unsupervised video for robust feature representation learning and then transfer to downstream tasks such as human action recognition, which deserves further exploration.
@@
Self-supervised learning has emerged as a promising approach for unsupervised learning. Its goal is to learn powerful feature representation by designing appropriate pretext tasks which automatically generate free annotations for unlabeled data and transfer the representations to various downstream tasks. This approach has achieved remarkable success in areas such as image recognition, object detection, and image segmentation. Besides the self-supervised learning on images [55, 60], there has been growing interest in exploring self-supervised learning for video representation with the increasing availability of large-scale unlabeled video data.
@@
Early video self-supervised learning approaches mainly focus on effectively exploiting contextual relationships within videos [58, 56]. Over the years, self-supervised methods have become more powerful and diverse can generally be categorized into two groups: discriminative classification / regression methods [94, 86] and generative prediction ones [29, 6, 4]. Discriminative methods focus on learning to distinguish between different samples, while generative methods aim to predict missing information. Recent efforts have been made to learn general video representations that can be applied to a range of downstream tasks. This includes video contrastive learning [94, 18, 69, 1], as well as cross-modal learning [115, 105], which learns from different modalities such as video, audio and text. Although these methods achieve promising results for various video based downstream tasks, they do not explicitly consider designing pretext tasks for action understanding, which may lead to sub-optimal solutions for action recognition. One interesting work, DCM [160] designs a two-stream network to separately learn action-related features, i.e., context and motion features, which achieves impressive results for action recognition. However, DCM fails to jointly consider the learning of context and motion features, lacking comprehensive view of these two features.
@@
In this work, we propose the Collaboratively Self-supervised Video Representation (CSVR) framework, which jointly learns dynamic motion and static context features for the downstream action recognition task. As illustrated in Figure 1, CSVR consists of three main representation learning branches, i.e., pose prediction branch, context matching branch, and a collaborative training branch named video generating branch. To enhance the performance of action recognition, we resort to the strong relationship between human pose and actions. Specifically, we employ a generative pose prediction branch based on a conditional generative adversarial network (CGAN) [83], which predicts future poses by taking current pose sequences as input. By excluding the potential motion noise from backgrounds, our method can effectively learn dynamic motion features highly related to action. Video key frames, known as the intra frame (I-frame) in video compression domain, contain valuable context information. To leverage the abundant pattern in I-frame for action recognition, we incorporate a discriminative context matching branch, which works by pulling the features of clips and I-frames from the same videos closer together, while simultaneously pushing apart pairs from different videos. Finally, we design the collaborative video generating branch to jointly optimize the dynamic and static features. The branch contains two video generating networks, one aims to reconstruct current video clips, another is trained for predicting future video clips. The video generation process requires both static context information and dynamic motion information, which provides a powerful collaborative learning objective, assisting the previous two branches in learning excellent static and dynamic features.
@@
Extensive experiments on three popular datasets with various backbones show that our CSVR achieves state-of-the-art performance for two downstream action understanding tasks, i.e., action recognition and video retrieval.
@@
1) We introduce a novel framework focusing on action recognition named Collaboratively Self-supervised Video Representation learning (CSVR), which simultaneously learns dynamic motion and static context features using three branches with discriminative and generative pretext tasks.
2) Considering the strong connections between human pose and actions, we design a generative pose prediction branch that predicts future poses from current pose sequences, enabling efficient and effective extraction of dynamic motion features.
3) Our carefully designed collaborative video generating branch enables joint optimization of dynamic motion and static context features, resulting in more comprehensive representations that benefit downstream action understanding tasks.
2 Related Work
As our method is primarily focused on self-supervised video representation learning for action recognition, in this section, we provide a brief overview of some typical works in the fields of video action recognition and self-supervised video representation learning. Additionally, we review some works on Generative Adversarial Networks (GAN) [84] as GAN-based generative branches play a crucial role in our method.
2.1 Video-Based Action Recognition
@@
2.2 Self-Supervised Video Representation Learning
@@
Although self-supervised video representation learning methods achieve impressive performance on various downstream tasks including action recognition, most of them are not specially designed for action-oriented representation learning. Differently, DCM [160] and P-HLVC [19] aims to learn generalizable features specific to action understanding. They simply extract motion features by contrastive learning and leverage diverse motion information, i.e., pose and motion vectors as self-supervised signals. Inspired by these approaches, we propose a Collaboratively Self-supervised Video Representation (CSVR) learning framework that simultaneously considers pose prediction and context matching as pretext tasks. By jointly optimizing the whole framework by video generating branch, CSVR provides a more comprehensive feature representation for downstream action recognition task.
2.3 Generative Adversarial Networks
@@
Generative adversarial networks (GAN) [84] usually consist of two encoder-decoder architecture networks: a generator producing fake data and a discriminator distinguishing between fake and real data. However, the inherent randomness of GAN output makes it difficult to control the generated data. To address this issue, Conditional GANs (CGAN) [83] are proposed for incorporating input conditions such as class labels to regulate the synthetic outputs. Recently, more efforts are devoted to video generation tasks. TGAN v2[77] is a video generation network that can efficiently generate realistic videos. Some works [163, 164] prove the feasibility of learning video representation from generative learning methods. We believe the process of generating diverse samples utilizing GANs can better capture human motion and learn representations which benefit the action understanding tasks. In CSVR, we employ 1D-CNN-based CGAN in pose prediction branch and R3D-based TGAN v2 in video generating branch. To enhance the stability and efficiency of our training process, we also adopt the gradient penalty from WGAN-GP [72] and the parameter update strategy from OSGANs [50]. These carefully-designed objectives help our method efficiently generate more accurate and diverse samples, thus improve the quality of learned dynamic motion representations.
3 Methodology
3.1 Overall Framework
@@
3.2 Generative Pose Prediction Branch
The generative pose prediction branch is conducted to learn dynamic motion features, which consists of two modules, i.e., a pose generator ($G_{p}$) and a pose discriminator ($D_{p}$), similar to the structure of CGAN [83]. Using the poses from current video clips as input, the objective of pose generator ($G_{p}$) is to generate plausible future poses. The feature from the last convolutional layers of encoder in $G_{p}$ serve as the source of the dynamic motion feature $f_{d}$. The structure of the pose prediction network is illustrated in the upper half of Figure 1.
Suppose $p_{t}\in\mathbb{R}^{2N}$ denotes the human pose captured in the $t$-th video frame, where $N$ is the number of skeleton joints. Given a randomly selected $T$-frame video clip, we apply OpenPose human skeleton detector [79] to estimate the current pose sequence $P_{in}=(p_{0},p_{1},\cdots,p_{T-1})\in\mathbb{R}^{2N}$ and the future pose sequence $P_{gt}=(p_{T},p_{T+1},...,p_{T+T^{\prime}-1})$ of length $T^{\prime}$ frames. The input to $G_{p}$ is a latent noise vector $z_{p}$ and the current pose sequence $P_{in}$. After applying a one-dimensional filter along the time direction, $G_{p}$ produces a sequence of poses for the next $T^{\prime}$ frames, denoted as $P_{gen}=(p^{\prime}_{T},p^{\prime}_{T+1},...,p^{\prime}_{T+T^{\prime}-1})$. During training, $G_{p}$ learns to generate pose sequences that match the target distribution, while $D_{p}$ aims to distinguish real pose sequences $P_{gt}$ from generated ones $P_{gen}$. The objective function is formulated as:
$\begin{split}\min_{G_{p}}&\max_{D_{p}}\mathcal{L}_{pose}=\\ &E_{P_{gt}}[\log D_{p}(P_{gt})]+E_{z}[\log(1-D_{p}(G_{p}(z_{p}|P_{in}))]+\\ &\frac{1}{T^{\prime}}\sum^{T+T^{\prime}-1}_{t=T}||p_{t}-p^{\prime}_{t}||_{2}^{2}+E_{p}[(\parallel\bigtriangledown D_{p}(P|P_{in})\parallel_{2}-1)^{2}]\end{split}$ (1)
@@ 3.2 Generative Pose Prediction Branch
3.3 Discriminative Context Matching Branch
@@
As depicted in the blue box of Figure 1, the V-Network (Video Network) and I-Network (I-frame network) are used to extract video and context features, respectively. We extract features $x_{i}\in\mathbb{R}^{C_{1}\times T_{1}\times H_{1}\times W_{1}}$ from a random video clip $i$ and feature $z_{i}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ from the corresponding I-frame of the same video, where $C_{1}$, $T_{1}$, $H_{1}$, and $W_{1}$ denote the channel, time, height, and width of video clips, respectively. We use global average pooling to obtain their global representations $x_{i}^{pool}\in\mathbb{R}^{C_{1}}$ and $z_{i}^{pool}\in\mathbb{R}^{C_{2}}$ We then apply two-layer MLP heads, $g^{V}$ and $g^{I}$, by taking the features of the video clip and I-frame as inputs, i.e., $x_{i}^{*}=g^{V}(x_{i}^{pool})\in\mathbb{R}^{C}$ and $z_{i}^{*}=g^{I}(z_{i}^{pool})\in\mathbb{R}^{C}$. Finally, the Multi-Instance-InfoNCE (MI-InfoNCE) loss [161] is applied to pull the video clip and I-frame feature pairs from the same video closer and push pairs from different video further away:
$\begin{split}\mathcal{L}_{context}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})}{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x^{*}_{i})}{\tau})}\end{split}$ (2)
@@
3.4 Collaborative Video Generating Branch

### v2

@@ Abstract
Considering the close connection between action recognition and human pose estimation, we design a Collaboratively Self-supervised Video Representation (CSVR) learning framework specific to action recognition by jointly factoring in generative pose prediction and discriminative context matching as pretext tasks. Specifically, our CSVR consists of three branches: a generative pose prediction branch, a discriminative context matching branch, and a video generating branch. Among them, the first one encodes dynamic motion feature by utilizing Conditional-GAN to predict the human poses of future frames, and the second branch extracts static context features by contrasting positive and negative video feature and I-frame feature pairs. The third branch is designed to generate both current and future video frames, for the purpose of collaboratively improving dynamic motion features and static context features. Extensive experiments demonstrate that our method achieves state-of-the-art performance on multiple popular video datasets.
Index Terms:
@@
Video representation learning, Self-supervised learning, Action recognition, Human pose prediction.
00footnotetext: Manuscript received 5 July 2024; revised 8 December 2024 and 1 January 2025; accepted 2 January 2025. Date of publication 16 January 2025; date of current version 15 January 2025. This work is partially supported by Natural Science Foundation of China (No. U2336213, 62176251) , Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680202), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No. SYG202325), and Youth Innovation Promotion Association CAS. The associate editor coordinating to the review of this manuscript amd approving it for publication was Prof. Zhen Lei. (Corresponding author: Shiguang Shan.) Jie Zhang, Zhifan Wan, Lanqing Hu and Shiguang Shan are with the Key Laboratory of AI Safety of CAS, Institute of Computing Technology (ICT), Chinese Academy of Sciences (CAS), Beijing 100190, China. They are also affiliated with the School of Computer Science and Technology, University of Chinese Academy of Sciences (UCAS), Beijing 100049, China (e-mail: zhangjie@ict.ac.cn; wanzhifan16@mails.ucas.ac.cn; lanqing.hu@vipl.ict.ac.cn; sgshan@ict.ac.cn). Stephen Lin is with Microsoft Research Asia, Beijing 100080, China (e-mail: stevelin@microsoft.com). Shuzhe Wu is with Beijing Huawei Digital Technologies Co., Ltd., No.3 Xinxi Road, Haidian District, Beijing 100095, China (e-mail: wushuzhe2@huawei.com).
I Introduction
@@
Human action recognition is a fundamental task in computer vision and has gained increasing attention in recent years. With a large number of labeled videos, the convolutional neural network (CNN) has achieved great success in human action recognition by learning powerful spatio-temporal representation features [32, 33, 81, 80]. However, the cost of collecting and tagging large amounts of videos is enormous. Fortunately, there are a large amount of unlabeled videos on the Internet. A common practice is to use a large amount of unsupervised video for robust feature representation learning and then transfer to downstream tasks such as human action recognition, which deserves further exploration.
@@
Self-supervised learning has emerged as a promising approach for unsupervised learning. Its goal is to learn powerful feature representation by designing appropriate pretext tasks which automatically generate free annotations for unlabeled data and transfer the representations to various downstream tasks. This approach has achieved remarkable success in areas such as image recognition, object detection, and image segmentation. In addition to self-supervised learning on images [55, 60], there has been a growing interest in exploring self-supervised learning for video representation with the increasing availability of large-scale unlabeled video data.
@@
Early self-supervised video learning approaches focus mainly on effectively exploiting contextual relationships within videos [58, 56, 12]. Over the years, self-supervised methods have become more powerful and can generally be categorized into two groups: discriminative classification / regression methods [94, 86] and generative prediction methods [168, 29, 6]. Discriminative methods focus on learning to distinguish between different samples, while generative methods aim to predict missing information. Recent efforts have been made to learn general video representations that can be applied to a range of downstream tasks. This includes video contrastive learning [94, 18, 69, 1, 65], as well as cross-modal learning [114, 105], which learns from different modalities such as video, audio, and text. Although these methods achieve promising results for various video based downstream tasks, they do not explicitly consider designing pretext tasks for action understanding, which may lead to sub-optimal solutions for action recognition. Considering the close connection between action recognition and human pose estimation, we design a self-supervised video representation learning framework specific to action recognition by explicitly inducing generative pose prediction as the pretext task. One interesting work, DCM [160] designs a two-stream network to separately learn action-related features, i.e., context and motion features, which achieves impressive results for action recognition. However, DCM fails to jointly consider the learning of context and motion features, lacking comprehensive view of these two features. Differently, we conduct collaboratively self-supervised video feature learning by jointly considering generative pose prediction and discriminative context matching as pretext tasks.
@@
In this work, we propose the Collaboratively Self-supervised Video Representation (CSVR) framework, which jointly learns dynamic motion and static context features for the downstream action recognition task. As illustrated in Figure 1, CSVR consists of three main representation learning branches, i.e.pose prediction branch, context matching branch, and a collaborative training branch named video generating branch. To enhance the performance of action recognition, we resort to the strong relationship between human pose and actions. Specifically, we employ a generative pose prediction branch based on a conditional generative adversarial network (CGAN) [83], which predicts future poses by taking current pose sequences as input. By excluding the potential motion noise from backgrounds, our method can effectively learn dynamic motion features highly related to action. Intra frame (I-frame) refers to a type of frame in video compression that contains complete image data and can be used as a reference point for decoding other frames. It is typically considered to to contain valuable context information. To leverage the abundant pattern in I-frame for action recognition, we incorporate a discriminative context matching branch, which works by pulling the features of clips and I-frames from the same videos closer together, while simultaneously pushing apart pairs from different videos. Finally, we design the collaborative video generating branch to jointly optimize the dynamic and static features. The branch contains two video generating networks, one aims to reconstruct current video clips, another is trained for predicting future video clips. The video generation process requires both static context information and dynamic motion information, which provides a powerful collaborative learning objective, assisting the previous two branches in learning excellent static and dynamic features.
@@
Extensive experiments on four popular datasets with various backbones show that our CSVR achieves state-of-the-art performance for two downstream action understanding tasks, i.e.action recognition and video retrieval.
@@
1) Considering the strong connections between human pose and actions, we introduce a novel framework focusing on action recognition named Collaboratively Self-supervised Video Representation learning (CSVR), which simultaneously learns dynamic motion and static context features using three branches with discriminative and generative pretext tasks.
2) Our carefully designed collaborative video generating branch enables joint optimization of dynamic motion and static context features, resulting in more comprehensive representations that benefit downstream action understanding tasks.
II Related Work
As our method is primarily focused on self-supervised video representation learning for action recognition, in this section, we provide a brief overview of some typical works in the fields of video action recognition and self-supervised video representation learning. Additionally, we review some generation methods as GAN-based [84] generative branches play a crucial role in our method.
II-A Video-Based Action Recognition
@@
II-B Self-Supervised Video Representation Learning
@@
Although self-supervised video representation learning methods achieve impressive performance on various downstream tasks including action recognition, most of them are not specially designed for action-oriented representation learning. Differently, DCM [160] and P-HLVC [19] aims to learn generalizable features specific to action understanding. They simply extract motion features by contrastive learning and leverage diverse motion information, i.e.pose and motion vectors as self-supervised signals. Inspired by these approaches, we propose a Collaboratively Self-supervised Video Representation (CSVR) learning framework that simultaneously considers pose prediction and context matching as pretext tasks. By jointly optimizing the whole framework by video generating branch, CSVR provides a more comprehensive feature representation for downstream action recognition task.
II-C Video Generation
@@
Over the past years, many efforts are devoted to video generation, which can be generally categorized into three groups: Generative Adversarial Networks (GAN) [84], Variational Autoencoders (VAE) [170] and Diffusion Models [171]. Generative adversarial network (GAN) [84] is initially used to generate high-fidelity samples that matched the data distribution. Some work[14, 168] also employed it as a pretext task for video representation learning. GANs usually consist of two encoder-decoder architecture networks: a generator producing fake data and a discriminator distinguishing between fake and real data. However, the inherent randomness of GAN output makes it difficult to control the generated data. To address this issue, Conditional GANs (CGAN) [83] are proposed for incorporating input conditions such as class labels to regulate the synthetic outputs. Recently, more efforts are devoted to video generation tasks. TGAN v2[77] is a video generation network that can efficiently generate realistic videos. Some works [14, 163, 164, 168] prove the feasibility of learning video representation from generative learning methods. VideoMoCo[14] is the first work to use GANs for video representation learning. It reconstructs the foreground and background of videos separately using two GANs and applies the encoder of generator to action recognition downstream tasks. TS-GAN[168] employs a spatio-temporal two-stream GAN network to reconstruct keyframes and optical flow information, leveraging the learned features in downstream tasks. Unfortunately, both approaches underperformed across multiple datasets. Although these methods using GANs can effectively reconstruct video dynamics, their features do not perform well in downstream classification tasks.
Variational Autoencoders (VAEs) [170] have become widely adopted in two-stage generative models, where the first stage compresses images into a lower-dimensional latent representation, and the second stage generates images from this latent space. In the context of video generation, 2D VAEs are often extended to 3D variants by incorporating 3D convolutional layers or temporal attention mechanisms [172, 173, 174, 185]. These approaches have demonstrated promising results in video generation. More recently, diffusion models have achieved remarkable success in both image [175, 176, 177, 178] and video generation [179, 180, 181, 182, 183, 184, 186]. Video diffusion models can be broadly categorized into pixel-space methods [179, 183] and latent-space methods [181, 182]. Compared to the pixel-space methods, latent-space methods offer significant efficiency advantages for video generation. However, they still incur high computational costs, which makes them challenging to apply in self-supervised video representation learning. Therefore, in our CSVR, we employ 1D-CNN-based CGAN in pose prediction branch and R3D-based TGAN v2 in video generating branch. To enhance the stability and efficiency of our training process, we also adopt the gradient penalty from WGAN-GP [72] and the parameter update strategy from OSGANs [50]. These carefully-designed objectives help our method efficiently generate more accurate and diverse samples, thus improve the quality of learned dynamic motion representations.
III Methodology
III-A Overall Framework
@@
III-B Generative Pose Prediction Branch
The generative pose prediction branch is conducted to learn dynamic motion features, which consists of two modules, i.e.a pose generator ($G_{p}$) and a pose discriminator ($D_{p}$), similar to the structure of CGAN [83]. Using the poses from current video clips as input, the objective of pose generator ($G_{p}$) is to generate plausible future poses. The feature from the last convolutional layers of encoder in $G_{p}$ serve as the source of the dynamic motion feature $f_{d}$. The structure of the pose prediction network is illustrated in the upper half of Figure 1.
Suppose $p_{t}\in\mathbb{R}^{2N}$ denotes the human pose captured in the $t$-th video frame, where $N$ is the number of skeleton joints. Given a randomly selected $T$-frame video clip, we apply OpenPose human skeleton detector [79] to estimate the current pose sequence $P_{in}=(p_{0},p_{1},\cdots,p_{T-1})\in\mathbb{R}^{1\times(T\times 2N)}$ and the future pose sequence $P_{gt}=(p_{T},p_{T+1},...,p_{T+T^{\prime}-1})$ of length $T^{\prime}$ frames. The input to $G_{p}$ is a latent noise vector $z_{p}$ and the current pose sequence $P_{in}$. After applying a one-dimensional filter along the time direction, $G_{p}$ produces a sequence of poses for the next $T^{\prime}$ frames, denoted as $P_{gen}=(p^{\prime}_{T},p^{\prime}_{T+1},...,p^{\prime}_{T+T^{\prime}-1})$. During training, $G_{p}$ learns to generate pose sequences that match the target distribution, while $D_{p}$ aims to distinguish real pose sequences $P_{gt}$ from generated ones $P_{gen}$. The objective function is formulated as:
$\begin{split}\min_{G_{p}}&\max_{D_{p}}\mathcal{L}_{pose}=\\ &\mathbb{E}_{p\sim P_{gt}}[\log D_{p}(P_{gt})]+\mathbb{E}_{z\sim p(z)}[\log(1-D_{p}(G_{p}(z_{p}|P_{in}))]+\\ &\frac{1}{T^{\prime}}\sum^{T+T^{\prime}-1}_{t=T}||p_{t}-p^{\prime}_{t}||_{2}^{2}+\mathbb{E}_{p}[(\parallel\bigtriangledown D_{p}(P|P_{in})\parallel_{2}-1)^{2}]\end{split}$ (1)
@@ III-B Generative Pose Prediction Branch
III-C Discriminative Context Matching Branch
@@
As depicted in the blue box of Figure 1, the V-Network (Video Network) and I-Network (I-frame network) are used to extract video and context features, respectively. We extract features $x_{i}\in\mathbb{R}^{C_{1}\times T_{1}\times H_{1}\times W_{1}}$ from a random video clip $i$ and feature $z_{i}\in\mathbb{R}^{C_{2}\times H_{2}\times W_{2}}$ from the corresponding I-frame of the same video, where $C_{1}$, $T_{1}$, $H_{1}$, and $W_{1}$ denote the channel, time, height, and width of video clips, respectively. We use global average pooling to obtain their global representations $x_{i}^{pool}\in\mathbb{R}^{C_{1}}$ and $z_{i}^{pool}\in\mathbb{R}^{C_{2}}$ We then apply two-layer MLP heads, $g^{V}$ and $g^{I}$, by taking the features of the video clip and I-frame as inputs, i.e.$x_{i}^{*}=g^{V}(x_{i}^{pool})\in\mathbb{R}^{C}$ and $z_{i}^{*}=g^{I}(z_{i}^{pool})\in\mathbb{R}^{C}$. Finally, the Multi-Instance-InfoNCE (MI-InfoNCE) loss [161] is applied to pull the video clip and I-frame feature pairs from the same video closer and push pairs from different video further away:
$\begin{split}&\mathcal{L}_{context}=\\ &-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})}{\sum\exp(\frac{\cos(z^{*}_{i},x^{*}_{i})}{\tau})+\sum_{k\neq i}^{B}\exp(\frac{\cos(z^{*}_{k},x^{*}_{i})}{\tau})}\end{split}$ (2)
@@
III-D Collaborative Video Generating Branch